OpenStudy (anonymous):

witch method is being used, subtraction or elimination? M=MP3 H=headphones 5x’s the cost of mp3 players + 8x’s the total of a set of headphones that = $840 5m+8h=840 3x’s the total cost of an mp3 + 4x’s the cost of a set of headphones = $480 3m+4h= 480 Equations 5m+8h=840 5m=840-8h M=(1/5)*(840-8h) 3m+4h=480 3*(1/5)*(840-8h) +4h=480 3*(840-8h)+20h=2400 2520-24h+20h=2400 -4h=-120 H=30 M=(1/5)*(840-8h) M=(1/5) *(840-8*(30)) M=(1/5)*(840-240) M=(1/5)*(600) M=120

OpenStudy (nottim):

That's a super long formula...

OpenStudy (nottim):

unless im reading that wrong. HEY. DONT TRIPLE POST

OpenStudy (anonymous):

I know and i got the answer I just dont know witch formaula i used, i just did it.

OpenStudy (anonymous):

M=MP3 H=headphones 5x’s the cost of mp3 players + 8x’s the total of a set of headphones that = $840 5m+8h=840 3x’s the total cost of an mp3 + 4x’s the cost of a set of headphones = $480 3m+4h= 480 Equations 5m+8h=840 5m=840-8h M=(1/5)*(840-8h) 3m+4h=480 3*(1/5)*(840-8h) +4h=480 3*(840-8h)+20h=2400 2520-24h+20h=2400 -4h=-120 H=30 M=(1/5)*(840-8h) M=(1/5) *(840-8*(30)) M=(1/5)*(840-240) M=(1/5)*(600) M=120

OpenStudy (nottim):

oh. that's your work.

OpenStudy (anonymous):

yes, but idk what elemination is, and i dont wanna say subtract if its also elemination

OpenStudy (anonymous):

whta.the.hay.

OpenStudy (nottim):

he showed his work in the question box.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

is it elemination or subtraction?!?!?!?!

OpenStudy (nottim):

give me a moment. IT's a bit confusing to read.

OpenStudy (nottim):

sorry, but I can't udnerstand what's going on.

OpenStudy (nottim):

i was going to say substitution, but that must be way off.

OpenStudy (nottim):

hey, you able to describe to me what's happening there?

OpenStudy (anonymous):

heres the question that i answered

OpenStudy (anonymous):

Translate the following into a system of equations, then solve it: A customer walks into an electronics store and buys five MP3 players and eight sets of headphones, paying $840. A second customer buys three MP3 players and four sets of headphones, and pays $480. How much does an MP3 player cost? How much does a set of headphones cost?

OpenStudy (anonymous):

and i just did it, i dont know what i used I JUST DID THE PROBLEM. M=MP3 H=headphones 5x’s the cost of mp3 players + 8x’s the total of a set of headphones that = $840 5m+8h=840 3x’s the total cost of an mp3 + 4x’s the cost of a set of headphones = $480 3m+4h= 480 Equations 5m+8h=840 5m=840-8h M=(1/5)*(840-8h) 3m+4h=480 3*(1/5)*(840-8h) +4h=480 3*(840-8h)+20h=2400 2520-24h+20h=2400 -4h=-120 H=30 M=(1/5)*(840-8h) M=(1/5) *(840-8*(30)) M=(1/5)*(840-240) M=(1/5)*(600) M=120

OpenStudy (nottim):

It looks like substitution.

OpenStudy (anonymous):

I agree!

OpenStudy (nottim):

So..I'm actually right, @Chlorophyll ?

OpenStudy (anonymous):

Absolutely, I'm tearing out my eyes to find it out finally :(

OpenStudy (anonymous):

5m=840-8h M=(1/5)*(840-8h)

OpenStudy (anonymous):

Still don't understand why people go super long way instead of 4 lines ???

OpenStudy (anonymous):

@NotTim you're greatly deserved medal for being pioneer of analyzing this "scary" post :)

OpenStudy (nottim):

Thanks. I do what I can to help on Os.

OpenStudy (anonymous):

I've never done this kind of math so i go the long way. MY BAD!